# G-Transformer: Counterfactual Outcome Prediction under Dynamic and Time-varying Treatment Regimes

### v2

@@ Abstract
In the context of medical decision making, counterfactual prediction enables clinicians to predict treatment outcomes of interest under alternative courses of therapeutic actions given observed patient history. Prior machine learning approaches for counterfactual predictions under time-varying treatments focus on static time-varying treatment regimes where treatments do not depend on previous covariate history. In this work, we present G-Transformer, a Transformer-based framework supporting g-computation for counterfactual prediction under dynamic and time-varying treatment strategies. G-Transfomer captures complex, long-range dependencies in time-varying covariates using a Transformer architecture. G-Transformer estimates the conditional distribution of relevant covariates given covariate and treatment history at each time point using an encoder architecture, then produces Monte Carlo estimates of counterfactual outcomes by simulating forward patient trajectories under treatment strategies of interest. We evaluate G-Transformer extensively using two simulated longitudinal datasets from mechanistic models, and a real-world sepsis ICU dataset from MIMIC-IV. G-Transformer outperforms both classical and state-of-the-art counterfactual prediction models in these settings. To the best of our knowledge, this is the first Transformer-based architecture for counterfactual outcome prediction under dynamic and time-varying treatment strategies. Code will be released upon publication of the paper.
@@ 1 Introduction
Clinicians often have to choose among multiple treatment options for their patients but do not have the ability to test every strategy before making a decision. Counterfactual prediction in medical decision-making involves the estimation of potential future trajectories of covariates of interest under alternative courses of action given observed history.
Treatment strategies of interest are typically time-varying, indicating decisions span multiple time points, and dynamic, implying that each treatment decision at a given time point is influenced by the preceding history up to that point. In selecting among competing dynamic treatment strategies, it is useful to obtain counterfactual predictions regarding a patient’s probability of experiencing adverse outcomes under each alternative strategy, based on their observed covariate history. Recent works presented deep learning approaches to estimate time-varying treatment effects (Lim et al., 2018; Bica et al., 2020a, b; Melnychuk et al., 2022) . However, most previous approaches focus on estimating counterfactual outcomes under static time-varying treatment strategies where treatments are not dependent on past covariate history.
In this work, we present G-Transformer, a transformer-based (Vaswani et al., 2017) framework for counterfactual prediction under dynamic and time-varying treatment strategies. G-Transformer supports g-computation, a causal inference technique for estimating treatment effects under dynamic treatment regimes (Robins, 1986, 1987). G-Transfomer estimates the conditional distribution of relevant covariates given covariate and treatment history at each time point using an encoder architecture, then produces Monte Carlo estimates of counterfactual outcomes by simulating forward patient trajectories under treatment strategies of interest. We evaluated G-Transformer extensively using multiple simulated datasets from two mechanistic models where ground-truths under the counterfactual strategies can be measured, and a real-world sepsis dataset to assess G-Transformer’s potential clinical utility in counterfactual predictions under alternative dynamic fluid administration regimes. Our contributions are the following:
  • G-Transformer architecture. We introduce a novel Transformer-based architecture that enables counterfactual predictions under dynamic and time-varying treatment strategies to provide estimates of individual or population-level treatment effects. We present a custom sequential training procedure for supporting g-computation that performs better than the seq2seq training typically used in Transformers.
@@
  • Evaluation under real-world ICU dataset. Using real-world data of a sepsis cohort from the MIMIC-IV database (Johnson et al., 2023b), we evaluated the performance of G-Transformer in predicting outcomes under observational treatment regimes, and demonstrated that G-Transformer can generate clinically meaningful counterfactual predictions under alternative dynamic fluids administration regimes.
@@
Our extensive experiments show strong empirical performance of G-Transformer under both static and dynamic time-varying treatment strategies over other state-of-the-art techniques. G-Transformer can facilitate treatment decision making at individual patient level by estimating patient trajectories under different treatment strategies of interest. At a population health level, G-Transformer can potentially be used to estimate population-level treatment effects to inform clinical guidelines and health policies. Although we focus on applications in healthcare in this proposed work, the G-Transformer architecture can be applied to other sequential decision making tasks that involve dynamic and time-varying interventions.
@@ 2 Related Work
Recent work by Wu et al. (2023) presented an alternating sequential model that used Transformers for clinical treatment outcome prediction, but their work did not support counterfactual prediction. Seedat et al. (2022) and Holt et al. (2024) focused on Continuous-Time Modeling, attempting to predict counterfactual treatment effects over irregular time lengths. They utilized neural controlled differential equations to model patient historical trajectories and attempted to make uncertainty estimates based on this model. T4 (Liu et al., 2023) introduced a LSTM-based neural network architecture for predicting the counterfactual treatment effects of sepsis. They also estimated the uncertainty associated with treatment effects.
@@ 3.1 Background: G-computation and Problem Setup
Assumption 3.1. Consistency asserts that the observed outcome aligns with the counterfactual outcome that would have occurred had the observed treatment been applied.
@@
G-computation algorithm requires the ability to simulate from joint conditional distributions $p(L_{t}|\bar{L}_{t-1},\bar{A}_{t-1})$ of the covariates given patient history at time $t$. These conditional distributions need to be estimated from data. Most implementations of g-computation use generalized linear regression models (GLMs) to estimate the conditional distributions of the covariates. These models often do not capture the complex temporal dependencies in the patient data. We propose the G-Transformer for this task.
@@ 3.2 G-Transformer Architecture
Figure 1: G-Transformer Architecture. G-Transformer uses two distinct encoders for categorical ($L^{ca}$) and continuous variables ($L^{co}$) respectively. The diagram illustrates (i) the training process through time steps T-1 and T, and (ii) the counterfactual simulation process starting at time T, with the counterfactual strategy initiated at the end of time T-1. Teacher forcing is used during training.
We utilized two Transformer encoders as the sequential model to separately learn hidden representations for continuous and categorical covariates in G-Transformer. We use the teacher forcing in the training process and g-computation at simulation process. The architecture is shown in Fig.1.
@@
In G-Transformer, we divide covariates into two groups: one for categorical variables $L_{t}^{ca}$, and the other for continuous variables $L_{t}^{co}$. Thus, $p(L_{t}|\bar{L}_{t-1},\bar{A}_{t-1})$ is defined as:
$p(L_{t}|\bar{L}_{t-1},\bar{A}_{t-1})=p(L^{ca}_{t}|\bar{L}_{t-1},\bar{A}_{t-1})\times p(L^{co}_{t}|L^{ca}_{t},\bar{L}_{t-1},\bar{A}_{t-1})$ (6)
Although assigning a customized model to each covariate may achieve better performance, considering the limitations of computational resources and the potential dependencies among variables, we decide to have only two Transformer encoder models to perform representation learning for categorical variables and continuous variables, respectively. At each time step $t$, we can compute represenetation of patient history $R_{t}=r_{t}(\bar{L}_{t},\bar{A}_{t};\theta)$, where $\theta$ denotes the learnable parameters from model $r_{t}$. We will introduce the specific content of model $r_{t}$ in the following section.
@@
Afterwards, we feed the combined representations of treatments and outcomes $h_{ca}^{e}$ into a Transformer encoder to obtain the final representation of past covariate history:
@@
The internal architecture of the Transformer encoder sublayers follows the original Transformer paper (Vaswani et al., 2017), including a multi-head self-attention layer, feed-forward network, and layer normalization mechanism. After obtaining $h_{ca}$, we map the hidden representation to the predicted future time step through a linear layer:
$\bar{L}_{t+1}^{ca}=Linear(ReLU(h_{ca}))$ (9)
@@
As with categorical variables, we use a Transformer encoder to obtain a hidden representation $h_{co}$ that incorporates the guiding variables. Then we utilize a linear layer to modify the hidden state to final output $\bar{L}_{t+1}^{co}$:
@@ 3.3 Training of G-Transformer
Since G-computation attempts to model the joint conditional probability distribution of the next time step $p(L_{t}|\bar{L}_{t-1},\bar{A}_{t-1})$ given the observable sequence, this approach is more aligned with the modeling methods of some RNN-like linear regression models. However, due to the Transformer’s nature as a seq-2-seq model, its multi-head self-attention mechanism endows it with powerful parallel computing capabilities. Therefore, the Transformer encoder is more inclined to model the probability distribution of a sequence, that is $p(\bar{L}_{2:t}|\bar{L}_{1:t-1},\bar{A}_{1:t-1})$. We tried two different training methods on our counterfactual prediction task, and the results are shown in the Appendix A.1. Furthermore, in real-world scenarios, most treatment variables appear in binary form, with few treatment variables presented in continuous form. Therefore, we believe that prioritizing modeling for categorical variables can provide more insights for the prediction of continuous variables. This approach also aligns more closely with the fundamental procedural form we use in the simulation process to calculate treatment variable $a_{m}=g_{m}(H_{m})$ in Algorithm. 2. After completing the entire prediction process, we use cross-entropy loss and MSE (Mean Squared Error) loss to optimize the prediction results for categorical and continuous variables, respectively. For categorical variables:
$\mathcal{L}_{ce}=-\frac{1}{N\times P_{ca}}\sum_{i}^{N}\sum_{c=1}^{P_{ca}}y_{ic}\;log(p_{ic};\theta)$ (13)
where $P_{ca}$ denotes the number of categorical variables, N denotes the number of paitents, and $\theta$ denotes the model parameters. In fact, we separately calculated the cross-entropy losses of different variables and computed their averages to account for the needs of multi-class variables. For continuous variables, we utilize the MSE loss:
$\mathcal{L}_{mse}=\frac{1}{N}\sum_{i=1}^{N}\sum_{t=1}^{K}||\bar{L}_{t,i}^{co}-\hat{L}_{t,i}^{co}||^{2}$ (14)
where $\hat{L}_{t,i}^{co}$ denotes the output of continuous encoder part of G-Transformer for patient $i$ in time $t$, $N$ denotes number of patients and $K$ denotes length of time. Algorithm 1 illustrates the overall training process of G-Transformer.
Input: Training set $\mathcal{D}$ data tuples: $\mathcal{D}=\{(\bar{L},\bar{A})\}$ ; the number of the prediction length $l$, Model parameters $\theta_{ca}$ and $\theta_{co}$; the number of the learning epochs $E$.
@@
       for $t=2$ to $l$ do
             Concatenate treatment variables and outcome variables $\bar{Y}_{1:t-1}=\{\bar{L}_{1:t-1},\bar{A}_{1:t-1}\}$.
             Use categorical encoder to compute the $\bar{Y}_{2:t}^{ca}=f_{ca}(\bar{Y}_{1:t-1};\theta_{ca})$.
             Keep the last time step of the predicted categorical variable $\bar{Y}_{t}^{ca}=\{\bar{L}_{t}^{ca},A_{t}^{ca}\}$.
             Use teacher forcing to process the input of continuous encoder $\bar{Y}_{1:t-1}^{te}=\{\bar{L}_{1:t-1},\bar{A}_{1:t-1},\bar{L}_{t}^{ca}\}$.
             Use continuous encoder to compute the $\bar{Y}_{2:t}^{co}=f_{co}(\bar{Y}_{1:t-1}^{te};\theta_{co})$.
             Keep the last time step of the predicted continuous variable $\bar{Y}_{t}^{co}=\{\bar{L}_{t}^{co},A_{t}^{co}\}$.
             Save the last time step $Y_{t}^{co},Y_{t}^{ca}$.
@@
      Concatenate the results of all predicted time steps $\bar{Y}_{2:t}^{ca}=\{{Y}_{2}^{ca},...,{Y}_{t}^{ca}\}$ and $\bar{Y}_{2:t}^{co}=\{{Y}_{2}^{co},...,{Y}_{t}^{co}\}$.
       Update model parameters $\theta_{ca}$ and $\theta_{co}$ by minimizing (13) and (14) with gradient descent
@@ 3.4 Simulation
After having a trained sequential model $f(R;\theta)$, we can obtain the conditional expectation estimate $\mathbb{E}[L_{t}|\bar{L}_{t-1},\{L_{t}^{i}\}_{i=1,...,p},\bar{A}_{t-1}]$ of all $p$ covariates at time step $t$ through Monte Carlo simulation. Without making parametric assumptions, we could simulate from:
@@
where $\epsilon_{t}^{p}$ is a noise which draws from the empirical distribution of the residuals $L^{p}_{t}-\hat{L}^{p}_{t}$ in a holdout set. If the variable is binary, then $\epsilon_{t}^{p}$ here is sampled from a Bernoulli distribution parameterized by empirical data. Afterward, we use Algorithm 2 for $K$ times to perform multi-timestep simulations. Figure 1 illustrates the simulation process. Different from the training process, since we cannot access to the ground truth, we sample from the learned conditional distribution when performing counterfactual predictions. For categorical variables, we use the parameters derived from the result $\hat{L}^{ca}$ of the categorical encoder to parameterize a categorical distribution (or a Bernoulli for binary indicator variables), and sample from the corresponding distribution. For the continuous variables, we added noise sampled from the empirical distribution collected from the validation set to the output part of the continuous encoder.
@@
We repeat the g-computation simulation Algorithm $M$ times. At the end of this process, we have $M$ simulated draws of the counterfactual outcome for each time $t=\{m,\ldots,K\}$. For each $t$, the empirical distribution of these draws constitutes a Monte-Carlo approximation of the counterfactual outcome distribution (2). The sample averages of the draws at each time $t$ are an estimate of the conditional expectations (1) and can serve as point predictions for $Y_{t}(\bar{A}_{m-1},\underline{g}_{m})$ in a patient with history $H_{m}$.

### v4

@@ Abstract
In the context of medical decision making, counterfactual prediction enables clinicians to predict treatment outcomes of interest under alternative courses of therapeutic actions given observed patient history. In this work, we present G-Transformer for counterfactual outcome prediction under dynamic and time-varying treatment strategies. Our approach leverages a Transformer architecture to capture complex, long-range dependencies in time-varying covariates while enabling g-computation, a causal inference method for estimating the effects of dynamic treatment regimes. Specifically, we use a Transformer-based encoder architecture to estimate the conditional distribution of relevant covariates given covariate and treatment history at each time point, then produces Monte Carlo estimates of counterfactual outcomes by simulating forward patient trajectories under treatment strategies of interest. We evaluate G-Transformer extensively using two simulated longitudinal datasets from mechanistic models, and a real-world sepsis ICU dataset from MIMIC-IV. G-Transformer outperforms both classical and state-of-the-art counterfactual prediction models in these settings. To the best of our knowledge, this is the first Transformer-based architecture that supports g-computation for counterfactual outcome prediction under dynamic and time-varying treatment strategies.
@@ 1 Introduction
Clinicians often have to choose among multiple treatment options for their patients but do not have the ability to test every strategy before making a decision. In selecting among competing dynamic treatment strategies, it is useful to obtain counterfactual predictions regarding a patient’s probability of experiencing adverse outcomes under each alternative strategy, based on their observed covariate history.
Counterfactual prediction in medical decision-making involves the estimation of potential future trajectories of covariates of interest under alternative courses of action given observed history. Treatment strategies of interest are typically time-varying, indicating decisions span multiple time points, and dynamic, implying that each treatment decision at a given time point is influenced by the preceding history up to that point. Recent works presented deep learning approaches to estimate time-varying treatment effects (Lim et al., 2018; Bica et al., 2020a, b; Melnychuk et al., 2022). However, most previous approaches focus on estimating counterfactual outcomes under static time-varying treatment strategies where treatments are not dependent on past covariate history.
In this work, we present G-Transformer, a transformer-based (Vaswani et al., 2017) framework for counterfactual prediction under dynamic and time-varying treatment strategies. G-Transformer supports g-computation, a causal inference technique for estimating treatment effects under dynamic treatment regimes (Robins, 1986, 1987). G-Transformer estimates the conditional distribution of relevant covariates given covariate and treatment history at each time point using an encoder architecture, then produces Monte Carlo estimates of counterfactual outcomes by simulating forward patient trajectories under treatment strategies of interest. We evaluated G-Transformer extensively using multiple simulated datasets from two mechanistic models where ground-truths under the counterfactual strategies can be measured, and a real-world sepsis dataset to assess G-Transformer’s potential clinical utility in counterfactual predictions under alternative dynamic fluid administration regimes. Our contributions are the following:
  • G-Transformer architecture. We introduce a novel Transformer-based architecture that enables counterfactual predictions under dynamic and time-varying treatment strategies to provide estimates of individual treatment effects. We present a custom sequential training procedure for supporting g-computation that performs better than the seq2seq training typically used in Transformers.
@@
  • Evaluation under real-world ICU dataset. Using real-world data of a sepsis cohort from the MIMIC-IV database (Johnson et al., 2023b; Goldberger et al., 2000), we evaluated the performance of G-Transformer in predicting outcomes under observational treatment regimes, and demonstrated G-Transformer’s potential clinical utility in generating counterfactual predictions under alternative dynamic fluids administration regimes.
@@
G-Transformer can facilitate treatment decision making by predicting time-varying treatment outcomes over time under alternative dynamic treatment strategies of interest. Our extensive experiments show strong empirical performance of G-Transformer under both static and dynamic time-varying treatment strategies over other state-of-the-art techniques. Although we focus on applications in healthcare in this proposed work, the G-Transformer architecture can be applied to other sequential decision making tasks that involve dynamic and time-varying interventions.
@@ 2 Related Work
Schulam and Saria (2017) proposes an implementation of continuous-time g-computation, focusing on static, time-varying treatment strategies. While their approach could potentially be extended to accommodate dynamic strategies, it is known that Gaussian Processes (GPs) become intractable for large datasets. Recent work by Wu et al. (2023) presented an alternating sequential model that used Transformers for clinical treatment outcome prediction, but their work did not support counterfactual prediction. Seedat et al. (2022) and Holt et al. (2024) focused on Continuous-Time Modeling, attempting to predict counterfactual treatment effects over irregular time lengths. They utilized neural controlled differential equations to model patient historical trajectories and attempted to make uncertainty estimates based on this model. T4 (Liu et al., 2023) introduced a LSTM-based neural network architecture for predicting the counterfactual treatment effects of sepsis. They also estimated the uncertainty associated with treatment effects.
@@ 3.1 Background: G-computation and Problem Setup
Assumption 3.1. Consistency asserts that the observed outcome is equal to the counterfactual outcome that would have occurred had the observed treatment been applied.
@@
G-computation algorithm requires the ability to simulate from joint conditional distributions $p(L_{t}|\bar{L}_{t-1},\bar{A}_{t-1})$ of the covariates given patient history.These conditional distributions need to be estimated from data. Most implementations of g-computation use generalized linear regression models (GLMs) to estimate the conditional distributions of the covariates. These models often do not capture the complex temporal dependencies in the patient data. We propose the G-Transformer for this task.
@@ 3.2 G-Transformer Architecture
Figure 1: G-Transformer Architecture. In this diagram, G-Transformer uses two distinct encoders for categorical ($L^{ca}$) and continuous variables ($L^{co}$) respectively. The diagram illustrates (i) the training process through time steps T-1 and T, and (ii) the counterfactual simulation process starting at time T, with the counterfactual strategy $gc$ initiated at the end of time T-1. Teacher forcing is used during training.
We utilized two Transformer encoders as the sequential model to separately learn hidden representations for continuous and categorical covariates in G-Transformer. We use the teacher forcing in the training process. Although we have presented our model architecture with two Transformer encoders—one for continuous variables and another for categorical variables— G-Transformer can be extended to support multiple encoders. For instance, a separate Transformer encoder can be used to model the conditional distribution of each individual covariate.
@@
In this work, unless otherwise noted, we divide covariates into two groups, one for the categorical variables $L_{t}^{ca}$, and the other for the continuous variables $L_{t}^{co}$. For simplicity, our description assumes categorical variables are simulated before continuous variables within each time step. However, our implementation does not impose any specific ordering requirements among the covariates.
We use two Transformer encoder models to perform representation learning for categorical variables and continuous variables respectively. At each time step $t$, we can compute representation of patient history $R_{t}=r_{t}(\bar{L}_{t},\bar{A}_{t};\theta)$, where $\theta$ denotes the learnable parameters from model $r_{t}$. We will introduce the specific content of model $r_{t}$ in the following section. In this work, the outcome $Y_{t}$ can be deemed as a covariate in the vector $L_{t+1}$. For example, the outcome of interest can be blood pressure in the next time step.
@@
Afterwards, we feed the combined representations of treatments and outcomes $h_{ca}^{r}$ into a Transformer encoder to obtain the final representation of past covariate history:
@@
The internal architecture of the Transformer encoder sublayers follows the original Transformer paper (Vaswani et al., 2017), including a multi-head self-attention layer, feed-forward network, and layer normalization mechanism. After obtaining $h_{ca}$, we map the hidden representation to the predicted future time step $\hat{L}^{ca}_{t+1}$ through a linear layer, followed by a softmax layer for getting probability $p^{ca}$ of categorical variables before loss calculation.
$\hat{L}_{t+1}^{ca}=ReLU(Linear(h_{ca}))$ (8)
@@
As with categorical variables, we use a Transformer encoder to obtain a hidden representation $h_{co}$ that incorporates the preceding variables. Then we utilize a linear layer to modify the hidden state to generate the final output $\hat{L}_{t+1}^{co}$:
@@ 3.3 Training of G-Transformer
Since g-computation models the joint conditional probability distribution of the next time step $p(L_{t}|\bar{L}_{t-1},\bar{A}_{t-1})$ given the observable sequence, it aligns more closely with the modeling methods used in RNN-like regression models. However, due to the Transformer’s nature as a seq-2-seq model, with its multi-head self-attention mechanism, it is typically used to generate an entire sequence at a time. We developed a customized training procedure that is aligned with the g-computation algorithm (see Algorithm 1). As a comparison, we present results from alternative training methods on our counterfactual prediction task, including the conventional Seq-2-Seq approach, and the results are shown in the Appendix A.1.
We use cross-entropy loss and MSE (Mean Squared Error) loss to optimize the prediction results for categorical and continuous variables, respectively. For categorical variables:
$\mathcal{L}_{ce}=-\frac{1}{N(K-m)D_{ca}}\sum_{i=1}^{N}\sum_{d=1}^{D_{ca}}\sum_{t=m}^{K}\sum_{c=1}^{C}L^{c,ca}_{idt}log(p_{idt}^{c,ca})$ (12)
where $D_{ca}$ denotes the number of categorical variables, $N$ denotes the number of paitents, $\theta$ denotes the model parameters, and $C$ denotes the number of classes in this categorical variable. We assume the class label of $L^{ca}_{d}$ is represented as a one-hot encoded vector $(L^{1}_{d},L^{2}_{d},...,L^{C}_{d})$ where $L^{c}_{d}=1$ for the correct class and 0 for all other classes. We separately calculated the cross-entropy losses of different variables and computed their averages to account for the needs of multi-class variables. For continuous variables, we utilize the MSE loss:
$\mathcal{L}_{mse}=\frac{1}{N(K-m)D_{co}}\sum_{i=1}^{N}\sum_{d=1}^{D_{co}}\sum_{t=m}^{K}||L_{t,i}^{d,co}-\hat{L}_{t,i}^{d,co}||^{2}$ (13)
where $\hat{L}_{t,i}^{co}$ denotes the output of continuous encoder part of G-Transformer for patient $i$ in time $t$, $N$ denotes number of patients, $K$ denotes length of time and $D_{co}$ denotes the number of continuous variables. Algorithm 1 illustrates the overall training process of G-Transformer.
Input: Training set $\mathcal{D}$ data tuples: $\mathcal{D}=\{(\bar{L},\bar{A})\}$ ; The number of the prediction length $K$; The number of the observed historical length $m$; Model parameters $\theta_{ca}$ and $\theta_{co}$; The number of the learning epochs $E$,.
@@
       for $t=m$ to $K$ do
             Use categorical encoder to compute the $\hat{L}_{t}^{ca}=f_{ca}(\bar{L}_{1:t-1},\bar{A}_{1:t-1};\theta_{ca})$.
             Use continuous encoder to compute the $\hat{L}_{t}^{co}=f_{co}(\{\bar{L}_{1:t-1},\bar{A}_{1:t-1},{L}_{t}^{ca}\};\theta_{co})$.
             Save the $\hat{L}_{t}^{co},\hat{L}_{t}^{ca}$ for loss calculation.
@@
      Concatenate the results of all predicted time steps $\hat{L}_{m:K}^{ca}=\{\hat{L}_{m}^{ca},...,\hat{L}_{K}^{ca}\}$ and $\hat{L}_{m:K}^{co}=\{\hat{L}_{m}^{co},...,\hat{L}_{K}^{co}\}$.
       Update model parameters $\theta_{ca}$ and $\theta_{co}$ by minimizing (12) and (13) with gradient descent.
@@ 3.4 Simulation
After having trained a sequential model $f(R;\theta)$, we can perform Monte-Carlo simulations, and sequentially simulate from $p(L_{t}^{p}|\bar{L}_{t-1},L_{t}^{0},\ldots,L_{t}^{p-1},\bar{A}_{t-1})$ as follows. If $L_{t}^{p}$ is multinomial, its conditional expectation defines its conditional density. If $L_{t}^{p}$ has a continuous density, we simulate from its conditional distribution as follows. Without making parametric assumptions, we simulate from:
@@
where $\epsilon_{t}^{p}$ is a draw from the empirical distribution of the residuals $L^{p}_{t}-\hat{L}^{p}_{t}$ in a holdout set. We use Algorithm 2 for each time $t=\{m,....,K\}$ to perform multi-timestep simulations. Figure 1 illustrates the simulation process. Different from the training process, since we do not have access to the ground truth, we sample from the learned conditional distribution when performing counterfactual predictions. For categorical variables, the predicted probabilities from $\hat{L}^{ca}$, produced by the categorical encoder, are used to parameterize a categorical distribution (or a Bernoulli for binary indicator variables). We then sample from the resulting distribution. For the continuous variables, we added noise sampled from the empirical distribution collected from the validation set to the output of the continuous encoder.
@@
We repeat the g-computation simulation (Algorithm 2) $M$ times, where $M$ is the number of Monte-Carlo simulations. At the end of this process, we have $M$ simulated draws of the counterfactual outcome for each time $t=\{m,\ldots,K\}$, where $m$ is beginning time step of simulation and $K$ is the end time step of simulation. For each $t$, the empirical distribution of these draws constitutes a Monte-Carlo approximation of the counterfactual outcome distribution (2). The sample averages of the draws at each time $t$ are an estimate of the conditional expectations (1) and can serve as point predictions for $Y_{t}(\bar{A}_{m-1},\underline{g}_{m})$ in a patient with history $H_{m}$.